Determine teh equation. Write the answer in standard for using only integers as the coefficients.
the line through (-2,0) with slope -3/4

Answer by stanbon(75887)

Write the answer in standard for using only integers as the coefficients.
the line through (-2,0) with slope -3/4
----------------
Form: y = mx + b
You know x,y, and m; solve for "b":
0 = (-3/4)*(-2) + b
0 = (3/2) + b
b = -3/2
----------------
EQuATION:
y = (-3/4)x -(3/2)
Multiplythru by 4 to get:
4y = -3x - 6
Rearrange:
3x+4y = -6
================
Cheers,
Stan H.